# Context Is Not Comprehension

### v2

@@ Abstract
The dominant evaluation of Large Language Models has centered on their ability to surface explicit facts from increasingly vast contexts. While today’s best models demonstrate near-perfect recall on these tasks, this apparent success masks a fundamental failure in multi-step computation when information is embedded in a narrative. We introduce Verbose ListOps (VLO), a novel benchmark designed to isolate this failure. VLO programmatically weaves deterministic, nested computations into coherent stories, forcing models to track and update internal state rather than simply locate explicit values. Our experiments show that leading LLMs, capable of solving the raw ListOps equations with near-perfect accuracy, collapse in performance on VLO at just 10k tokens. The VLO framework is extensible to any verifiable reasoning task, providing a critical tool to move beyond simply expanding context windows and begin building models with the robust, stateful comprehension required for complex knowledge work.
@@ 1 Introduction
Despite boasting million-token context windows, today’s leading Large Language Models (LLMs) can fail at reasoning tasks a human finds trivial: tracking a multi-step argument buried within a distracting narrative. This ability—to filter irrelevant information and track intermediate conclusions for later synthesis—is the foundation of high-value knowledge work, from a lawyer interpreting intepreting clauses to a sales manager inferring customer intent from chatty transcripts. Yet this core skill remains a fundamental limitation as existing benchmarks are ill-equipped to measure it. To bridge this gap, we introduce Verbose ListOps (VLO), a novel benchmark that isolates this ability in a controlled setting, challenging models to reason, not just locate information.
@@
VLO is specifically designed to address the limitations of prior evaluation methods. Where benchmarks for sequence comprehension (An et al., 2023; Zhang et al., 2024) or needle-in-a-haystack factual recall (Kamradt, 2023; Li et al., 2024b) test factual extraction, VLO embeds deterministic, nested ListOps computations (Nangia and Bowman, 2018; Tay et al., 2021) inside lengthy, coherent stories. Where synthetic datasets can leak intermediate solutions that models merely recall (Fodor, 2025), VLO enforces a strict protocol that withholds every intermediate result, forcing models to compute and maintain values in ’working memory’. And while human-annotated datasets offer realism, they are too coarse-grained to disentangle narrative interference from reasoning difficulty (Wang et al., 2024; Shaham et al., 2022; Bowman and Dahl, 2021); VLO’s agentic generation provides orthogonal controls for both. This setup probes an LLM’s ability to maintain computational ’state’ and follow algorithmic logic under narrative camouflage—coherent, semantically-related distractors that obscure a task—prerequisites for text understanding comparable to human performance.
@@
In this paper, we deliver three key contributions:
  1. 1.
    A benchmark for narrative-embedded deterministic reasoning. We weave ListOps computations into coherent fictional stories and withhold the intermediate results of all the nested problems, forcing models to compute, store, and recall values internally.
  2. 2.
    Orthogonal control of context length and reasoning complexity. We provide independent parameters for narrative length and reasoning complexity, enabling systemic exploration of scaling behavior and failure modes.
  3. 3.
    An extensible generalizable generation framework. We open-source a fully programmatic agentic pipeline where "author" and "critic" LLMs collaborate to embed any deterministically verifiable task; from numerical operations to symbolic and logical reasoning, into a coherent narrative. This methodology yields reliable, scalable data and provides a foundation for testing a wide range of complex reasoning beyond the scope of this initial work.
2 Related Work
@@ 2.1 Long-Context Benchmarks
Early long-context evaluations typically adapted standard NLP benchmarks (Shaham et al., 2023; An et al., 2023; Bai et al., 2024). These early benchmarks often had contexts shorter than the maximum capabilities of current LLMs (Gemini Team, 2024; OpenAI, 2025a; Anthropic, 2025) and inadequately differentiated complex reasoning tasks from simpler simpler tasks based on factual recall tasks or failed to analyze the impact of distractors sufficiently (Muhlgay et al., 2023).
@@ 3.1 Formal Task Definition
$T=\textsc{max}(\textsc{sum}(2,1),\,4),$
then one computes $\textsc{sum}(2,1)=3$ and subsequently $\textsc{max}(3,4)=4$.
@@
with intermediate values appearing only as anchors ${a_{i}}$. Success is $f(X)$ matching the ground-truth integer.
@@
Consider the above ListOps problem again. The first operation is
$\textsc{sum}(2,1)\;\to\;a_{1},\quad a_{1}=3$
This intermediate value (3) is tracked by anchor $a_{1}$. Example narrative segment:
@@ 3.3 Generation Pipeline
Figure 2 outlines how the VLO construction pipeline ensures adherence to constraints C1–C3 through:
@@
    AST Sampling (Programmatic). Sample a ListOps AST $T$ of depth $d$ and branching factor $b$, and compute its root value $v(T)$ by deterministic post-order evaluation.
@@
    Iterative Operator Narrative Construction. For each ListOps node, build a narrative segment via:
@@
      Author Agent (LLM). The Author LLM drafts a segment for the current operator $\omega\in\mathcal{O}$ and its child anchors $a_{i}$ (if applicable), weaving it into the story (C2) while avoiding numeric mention of the current result (C1) and respecting parameters (C3).
@@
      Critic Agent (LLM). The Critic LLM reviews the draft, checking:
@@
      Static Validation (Programmatic). A script scans the finalized segment to verify:
@@
  3. 3.
    Holistic Narrative Validation (LLM). After concatenating all segments into the full narrative, validator.py orchestrates a review of the entire sample:
    • Comprehensive LLM validation: A larger state-of-the-art Validator LLM, aided by extensive prompting and few-shot examples, reads the entire story. It performs a step-by-step evaluation of the narrative against the original AST $T$ to detect subtle numeric leaks (e.g., spelled-out numbers for intermediate results), verify correct representation of operations and inputs (including conceptual anchors), check for missing or duplicate operators, ensure overall coherence (C2), and confirm that its own evaluation of the original AST $T$ matches the provided ground truth value $v(T)$.
    Violations identified by the Validator LLM cause the sample to be discarded and resampled.
Figure 2: The Verbose ListOps (VLO) agentic generation and validation pipeline, employing Author, Critic, and Validator LLMs alongside programmatic checks to ensure sample validity.
@@ 3.7 Results
Model ListOps VLO-10k OAI MRCR-8k 8-Pin 95% CI (ListOps) 95% CI (VLO)
@@
Table 2: Accuracy on VLO-10k versus its corresponding bare ListOps expressions, based on evaluation of 1,000 samples per model. Of particular interest is DeepSeek V3’s plunge in performance from 93.7% to 25.1%. OpenAI MRCR results (95% CI) sourced from ContextArena (Uzar, 2025).
@@ 3.8 Analysis and Limitations
The VLO benchmark’s core challenge is its use of narrative camouflage to obscure a nested deterministic task, creating a information processing burden that exposes fundamental architectural differences between models. The results in Table 2 reveal two critical takeaways. First, models with explicit reasoning scaffolds (e.g., Gemini 2.5, o4-Mini) perform better on VLO than those without. Second, the stark contrast between DeepSeek-V3’s near-perfect score on raw ListOps and its collapse on VLO demonstrates while scratchpad-like mechanisms are not required for nested reasoning, they are crucial for filtering noise and protecting the reasoning process from distractors.
This architectural divergence explains the performance trends. DeepSeek-V3’s design, while highly efficient, creates specific vulnerabilities to VLO’s challenge. Its architecture is optimized for performance on standard benchmarks through two key strategies that, we hypothesize, compromise its task resiliance on VLO. First, its auxiliary-loss-free load balancing for its Mixture-of-Experts (MoE) architecture encourages over-specialisation, where the gating network routes inputs to a "narrative" expert that fails to pattern-match the embedded logic. Second, its use of Multi-Token Prediction (MTP) encourages the model to "pre-plan" its output based on narrative flow, reinforcing the very heuristic processing that VLO penalizes (DeepSeek-AI et al., 2025).
@@
  • Generator Model Bias: VLO-10k was generated with Gemini 2.5.111During development we found only Gemini 2.5 and GPT-4.5 could reliably generate VLO, with GPT-4.5 pricing being cost-prohibitive (\$75USD/1M input tokens, \$150USD/1M output tokens (OpenAI, 2025b)). This introduces potential generator model bias, where evaluated models from the same family may exhibit inflated performance due to stylistic or implicit knowledge alignment. Future work should investigate this and explore mitigation strategies, such as employing a more diverse set of future generator models or incorporating bias-neutralization techniques (Yuan et al., 2025).
@@ 4 Discussion
Concept-oriented Large Concept Models (LCMs),222’LCMs’ here do not refer to Latent Consistency Models (Luo et al., 2023) used for image synthesis. which operate on higher-level semantic units rather than individual tokens (Meta LCM Team et al., 2024) offer a promising direction. For VLO, an LCM treats each reasoning step or conceptual reference to an intermediate result (our "narrative anchors") as distinct conceptual units. This aligns more closely with how humans might track such information and could make the model more resistant to narrative distractors when trying to maintain the integrity of a numerical value associated with a concept. In performing autoregressive prediction in a concept embedding space, rather than a token space, LCMs could facilitate more stable tracking of multi-step computations. Such concept embeddings can be a robust and less diffuse carrier of the numerical value of an intermediate result than distributed token activations.
@@ E.4 Main Beat Generation Prompt (Illustrative Core Rules)
  - **Action (Specific to Op, e.g., SUM):** Characters combine/tally ALL inputs...

### v4

@@ Abstract
The dominant way of judging Large Language Models (LLMs) has been to ask how well they can recall explicit facts from very long inputs. While today’s best models achieve near-perfect recall, this masks a harder skill: performing multi-step reasoning and tracking intermediate state that never appears verbatim. We introduce Verbose ListOps (VLO), a benchmark that embeds deterministic ListOps computations inside narrative camouflage and, crucially, allows step-level evaluation of every intermediate result. Experiments show that models which solve raw ListOps with $\approx 100\%$accuracy collapse on VLO after only 10 k tokens. By exposing where a model’s reasoning chain first diverges, VLO moves assessment beyond sheer context length and toward genuine comprehension. VLO’s generation pipeline is task-agnostic: it can weave any deterministically verifiable reasoning schema—arithmetic, symbolic, abductive, inductive or defeasible—into narrative form. This makes VLO a reusable test-bed for the next wave of reasoning-centric model designs, not merely those with step-explicit scaffolds.
@@ 1 Introduction
Despite million-token context windows, today’s LLMs still stumble on tasks that a human finds trivial: following a multi-step argument buried inside a distracting story (Liu et al., 2024; Shi et al., 2023; An et al., 2023). This ability—to ignore irrelevant detail, maintain intermediate conclusions, and synthesise them later—is the backbone of knowledge work, from a lawyer interpreting clauses to a sales manager inferring intent from chatty transcripts. Yet benchmarks rarely measure it. To bridge this gap, we introduce Verbose ListOps (VLO), a novel benchmark that isolates this ability in a controlled setting, challenging models to reason, not just locate information.
Existing long-context suites chiefly test (i) factual recall in "needle-in-a-haystack" settings (Kamradt, 2023; Li et al., 2024b) or (ii) shallow reading comprehension (An et al., 2023; Zhang et al., 2024). They do not control reasoning depth, nor do they hide intermediate answers, so models can succeed through pattern matching rather than computation.
Verbose ListOps (VLO) fills this gap. ListOps is a synthetic benchmark where models perform nested arithmetic operations on sequences of integers following a tree structure—for example, [MAX 3 7 [MIN 1 9 5] 2] requires computing the inner MIN first, then finding the maximum among all values. We convert each ListOps node into a short story fragment, hide every intermediate value behind a ’narrative anchor’, and add semantically related distractors. The resulting benchmark forces a model to compute, cache, and reuse internal state.
@@
In this paper, we deliver three key contributions:
  1. 1.
    A broadly-applicable, agentic generation framework. We open-source a "World-Author–Critic–Validator" pipeline that can embed any deterministically-verifiable reasoning graph—not only arithmetic ListOps but also symbolic, abductive, inductive, or defeasible logic (see Appendix A).
  2. 2.
    Verbose ListOps (VLO): a first instantiation of the framework. Using this pipeline we build VLO, a benchmark that narrativises nested ListOps computations. VLO is the first long-context dataset that (i) withholds all intermediate results, (ii) controls context length and reasoning depth orthogonally, and (iii) ships exact intermediate-step labels, enabling systematic failure-mode analysis.
  3. 3.
    Orthogonal control knobs for context and reasoning complexity. Tree depth, operator arity, narrative length, and distraction density are exposed as independent hyper-parameters, allowing researchers to stress-test scaling behaviour along a single axis while holding all others fixed—something prior benchmark suites cannot do.
  4. 4.
    A diagnostic evaluation toolkit and empirical study. We release scripts that check correctness at every anchor step, plot error-propagation curves, and compare prompting regimes. Applying this toolkit to ten prominent LLMs uncovers a consistent blind spot: models that excel at raw ListOps or factual recall often diverge at the very first hidden step when the same logic is wrapped in narrative camouflage.
Together, the framework plus toolkit turn VLO from a one-off dataset into a reusable test-bed for any future long-context reasoning architecture.
2 Related Work
@@ 2.1 Long-Context Benchmarks
Early long-context evaluations typically adapted standard NLP benchmarks (Shaham et al., 2023; An et al., 2023; Bai et al., 2024). These early benchmarks often had contexts shorter than the maximum capabilities of current LLMs (Gemini Team, 2024; OpenAI, 2025a; Anthropic, 2025) and inadequately differentiated complex reasoning tasks from simpler tasks based on factual recall tasks or failed to analyze the impact of distractors sufficiently (Muhlgay et al., 2023).
@@ 2.2 The Emergence of Reasoning alongside Long Context
As VLO exposes every intermediate state, it is also a natural playground for Process-Reward Models (PRMs) that score reasoning chains rather than final outputs. We expect the dataset to serve both as a training signal and as a forensic tool for such chain-inspection methods.
@@ 3.1 Formal Task Definition
$T=\texttt{MAX}(\texttt{SUM}(2,1),\,4),$
then one computes $\texttt{SUM}(2,1)=3$ and subsequently $\texttt{MAX}(3,4)=4$.
@@
…where v(T) denotes the scalar value at the root of the abstract-syntax tree (i.e. the final numerical result). Success is $f(X)$ matching the ground-truth integer.
@@
Consider the above ListOps problem again. Intermediate values appear only as anchors ${a_{i}}$. The first operation is
$\texttt{SUM}(2,1)\;\to\;a_{1},\quad a_{1}=3$
In VLO, this operation would be narrativized as the following:
@@ 3.3 Generation Pipeline
Figure 2 outlines how the VLO construction pipeline ensures adherence to constraints C1–C3 through the following five-stage process (implementation details and prompts are provided in Appendices D and E):
@@
    AST Sampling. Sample a ListOps AST $T$ of depth $d$ and branching factor $b$, and compute its root value $v(T)$ by deterministic post-order evaluation.
@@
    World Generation. Generate fictional world metadata (genre, setting, characters, and primary objects) using an LLM to establish thematic coherence for narrative embedding (constraint C2). This creates the contextual foundation for embedding ListOps operations into a cohesive story. Narrative anchors—conceptual names like "Daily Cell Intake"—are generated to reference intermediate results without revealing numerical values (constraint C1). Detailed parameters are in Appendix C.
  3. 3.
    Distraction Generation. Generate semantically relevant but computation-irrelevant content between ListOps operations. This expands context length while forcing models to filter computational information from extraneous narrative detail, using the same world context to maintain coherence without introducing interfering numerical information that can interfere with the underlying reasoning task. The token watcher monitors narrative length throughout generation, dynamically allocating the remaining token budget after ListOps narratives.
  4. 4.
    Operator Narrative Construction. For each ListOps node, build a narrative segment via:
@@
      Author Agent. The Author LLM drafts a segment for the current operator $\omega\in\mathcal{O}$ and its child anchors $a_{i}$ (if applicable), weaving it into the story (C2) while avoiding numeric mention of the current result (C1) and respecting parameters (C3).
@@
      Critic Agent. The Critic LLM reviews the draft, checking:
@@
      Programmatic Static Validation. A script scans the finalized segment to verify:
@@
  5. 5.
    Comprehensive LLM Validation: A larger Validator LLM, aided by extensive prompting and few-shot examples, performs step-by-step evaluation against the original AST $T$ to detect numeric leaks, verify correct representation of operations and inputs, check for missing/duplicate operators, ensure coherence (C2), and confirm ground truth value $v(T)$. Violations identified by the Validator LLM cause the sample to be discarded and resampled.
Figure 2: The Verbose ListOps (VLO) agentic generation and validation pipeline. The process begins with AST sampling and world generation, then iteratively constructs narrative segments using Author-Critic LLM pairs with programmatic validation, incorporates distraction content, and concludes with holistic validation to ensure sample validity and adherence to constraints C1-C3.
@@ 3.7 Results
Model ListOps VLO-10k OAI MRCR-8k 8-Pin [^1] OpenAI’s Multi-Reference Context Recall (MRCR) benchmark: 8-passage subset measuring pure retrieval accuracy under 8k tokens without multi-step reasoning. Provides baseline for comparing retrieval vs. VLO’s narrative-embedded reasoning. Results from ContextArena (Uzar, 2025). 95% CI (ListOps) 95% CI (VLO)
@@
Table 2: Accuracy on VLO-10k versus its corresponding bare ListOps expressions, based on evaluation of 1,000 samples per model. Note DeepSeek V3’s drop from 93.7% to 25.1%.
@@ 3.8 Analysis and Limitations
VLO challenges models to find, solve, track, and synthesize nested reasoning problems amongst narrative camouflage to answer a deterministic question, creating a information processing burden that exposes fundamental architectural differences between models. The results in Table 2 reveal two critical takeaways. First, models with explicit reasoning scaffolds (e.g., Gemini 2.5, o4-Mini) perform better on VLO than those without. Second, the stark contrast between DeepSeek-V3’s near-perfect score on raw ListOps and its collapse on VLO demonstrates while scratchpad-like mechanisms are not required for nested reasoning, they are crucial for filtering noise and protecting the reasoning process from distractors.
This architectural divergence explains the performance trends. DeepSeek-V3’s design, while highly efficient, creates specific vulnerabilities to VLO’s challenge. Its architecture is optimized for performance on standard benchmarks through two key strategies that, we hypothesize, compromise its task resilience on VLO. First, its auxiliary-loss-free load balancing for its Mixture-of-Experts (MoE) architecture encourages over-specialisation, where the gating network routes inputs to a "narrative" expert that fails to pattern-match the embedded logic. Second, its use of Multi-Token Prediction (MTP) encourages the model to "pre-plan" its output based on narrative flow, reinforcing the very heuristic processing that VLO penalizes (DeepSeek-AI et al., 2025).
@@
  • Generator Model Bias: VLO-10k was generated with Gemini 2.5.222During development we found only Gemini 2.5 and GPT-4.5 could reliably generate VLO, with GPT-4.5 pricing being cost-prohibitive (\$75USD/M input tokens, \$150USD/M output tokens (OpenAI, 2025b)). This introduces potential generator model bias, where evaluated models from the same family may exhibit inflated performance due to stylistic or implicit knowledge alignment. Future work should investigate this and explore mitigation strategies, such as employing a more diverse set of future generator models or incorporating bias-neutralization techniques (Yuan et al., 2025).
@@ 4 Discussion
Concept-oriented Large Concept Models (LCMs),333’LCMs’ here do not refer to Latent Consistency Models (Luo et al., 2023) used for image synthesis. which operate on higher-level semantic units rather than individual tokens (Meta LCM Team et al., 2024) offer a promising direction. For VLO, an LCM treats each reasoning step or conceptual reference to an intermediate result (our "narrative anchors") as distinct conceptual units. This aligns more closely with how humans might track such information and could make the model more resistant to narrative distractors when trying to maintain the integrity of a numerical value associated with a concept. In performing autoregressive prediction in a concept embedding space, rather than a token space, LCMs could facilitate more stable tracking of multi-step computations. Such concept embeddings can be a robust and less diffuse carrier of the numerical value of an intermediate result than distributed token activations.
@@ E.4 Main Beat Generation Prompt (Illustrative Core Rules)
[…]